1,2,9,28,65... which number is next of this series?
Answer by rothauserc(4718)

the nth term Xn = n^3 + 1
:
we are given
:
1, 2, 9, 28, 65, ?
:
******************
X5 = 5^3 + 1 = 126
******************
: